OpenStudy (anonymous):

Find the slope of the line containing the points (4, -7) and (-3, 3).

OpenStudy (anonymous):

For this problem we can use the slope formula, the formula states: m=y2-y1/x2-x. Where m is the slope. Using the coordinates that you provided we get: 3-(-7)/-3-4. After simplifying we get: 3+7/-3-4= -10/7. This is our slope.